# V2X-Radar: A Multi-modal Dataset with 4D Radar for Cooperative Perception

### v2

@@ Abstract
Modern autonomous vehicle perception systems often struggle with occlusions and limited perception range. Previous studies have demonstrated the effectiveness of cooperative perception in extending the perception range and overcoming occlusions, thereby enhancing the safety of autonomous driving. In recent years, a series of cooperative perception datasets have emerged; however, these datasets primarily focus on cameras and LiDAR, neglecting 4D Radar—a sensor used in single-vehicle autonomous driving to provide robust perception in adverse weather conditions. In this paper, to bridge the gap created by the absence of 4D Radar datasets in cooperative perception, we present V2X-Radar, the first large-scale, real-world multi-modal dataset featuring 4D Radar. V2X-Radar dataset is collected using a connected vehicle platform and an intelligent roadside unit equipped with 4D Radar, LiDAR, and multi-view cameras. The collected data encompasses sunny and rainy weather conditions, spanning daytime, dusk, and nighttime, as well as various typical challenging scenarios. The dataset consists of 20K LiDAR frames, 40K camera images, and 20K 4D Radar data, including 350K annotated boxes across five categories. To support various research domains, we have established V2X-Radar-C for cooperative perception, V2X-Radar-I for roadside perception, and V2X-Radar-V for single-vehicle perception. Furthermore, we provide comprehensive benchmarks across these three sub-datasets. We will release all datasets [^1] http://openmpd.com/column/V2X-Radar and benchmark codebase [^2] https://github.com/yanglei18/V2X-Radar.
Figure 1: A data frame sampled from the V2X-Radar dataset. Each sample includes data from three sensors: (1) dense point clouds (gray points) from roadside and vehicle-side LiDAR; (2) sparse point clouds (green and blue points) with Doppler information from the roadside and vehicle-mounted 4D Radar; (3) RGB images (top row) from the vehicle’s camera and multi-view roadside cameras. All sensors are temporally and spatially synchronized. Each data frame is manually annotated with 3D boxes across five categories.
Table 1: Comparisons of our proposed V2X-Radar with existing autonomous driving and V2X datasets. C: Camera, L: Lidar, R: Radar, V2V: vehicle-to-vehicle, V2I: Vehicle-to-infrastructure.
Dataset Reference Location Sensor V2X Real 4D Radar Adverse Weather Day&Night #LiDAR #Images #4D Radar #3D Box
OpenV2V [37] ICRA’22 Carla L&C V2V 11K 44K 0 233K
V2X-Set [36] ECCV’22 Carla L&C V2V&I 11K 44K 0 233K
DAIR-V2X [48] CVPR’22 China L&C V2I 39K 39K 0 464K
V2X-Seq [49] CVPR’23 China L&C V2I 39K 39K 0 464K
V2V4Real [39] CVPR’23 USA L&C V2V 20K 40K 0 240K
RCooper [7] CVPR’24 China L&C I2I 30K 50K 0 310K
TUMTraf-V2X[55] CVPR’24 DE L&C V2I 2K 5K 0 29.38K
[…]
@@ 1 Introduction
Perception is critical in autonomous driving. While many single-vehicle perception methods [29, 28, 16, 26, 54, 43, 23, 23, 22, 25, 14] have emerged, substantial safety challenges persist due to occlusions and limited perception ranges. These issues occur because vehicles can only view their surroundings from a single perspective, resulting in an incomplete understanding of the scenario. This limitation hinders autonomous vehicles from achieving safe navigation and making optimal decisions. To address this challenge, recent studies have explored cooperative perception, where ego vehicles extend their perception range and overcome occlusions with assistance from other vehicles or roadside perception [42, 41, 44, 45, 31].
@@
To address the gap in the 4D Radar dataset for cooperative perception and to facilitate related studies, we present V2X-Radar, the first large-scale real-world multi-modal dataset featuring 4D Radar. This dataset includes a diverse range of scenarios, covering various weather conditions such as sunny, rainy, and snowy environments, and spans different times of the day, including daytime, dusk, and nighttime. The data was collected using a connected vehicle platform alongside an intelligent roadside unit, both equipped with 4D Radar, LiDAR, and multi-view cameras (Fig. 1). From over 15 hours of driving logs, we meticulously selected 50 representative scenarios for the final dataset. It comprises 20K LiDAR frames, 40K camera images, and 20K 4D Radar data, featuring 350K annotated bounding boxes across five object categories. To support a variety of research areas, V2X-Radar is further divided into three specialised sub-datasets: V2X-Radar-C for cooperative perception, V2X-Radar-I for roadside perception, and V2X-Radar-V for single-vehicle perception. Additionally, we provide comprehensive benchmarks of recent perception algorithms across these three sub-datasets. In comparison to existing real-world cooperative datasets, our V2X-Radar demonstrates two key strengths: (1) More modalities: The proposed V2X-Radar dataset includes three types of sensors: LiDAR, Camera, and 4D Radar, enabling further exploration into 4D Radar-related cooperative perception research. (2) Diverse scenarios: Our data collection covers diverse weather conditions and times of day, focusing on complex intersections that pose challenges for single-vehicle autonomous driving. These scenarios feature obstructed blind spots that impact vehicle safety, providing varied corner cases for cooperative perception research.
@@
  • Comprehensive benchmarks of recent perception algorithms are conducted across cooperative perception on V2X-Radar-C, roadside perception on V2X-Radar-I, and vehicle-side perception on V2X-Radar-V. We will release the entire V2XS-Radar dataset and benchmark codebase.
@@ 2.1 Autonomous Driving Datasets
Public datasets [1, 27, 37, 36, 12, 49, 48, 39, 32, 19, 52, 53] have significantly accelerated the progress of autonomous driving in recent years. Single-vehicle datasets, such as KITTI [6], NuScenes [1], and Waymo [27], have notably furthered the development of single-vehicle perception. In terms of cooperative perception datasets, OPV2V [37] is the first dataset in this area, gathering data through co-simulation using CARLA [3] and OpenCDA [34, 38]. V2XSet [36] and V2X-Sim [12] explore V2X perception by employing synthesized data from the CARLA simulator [3]. Unlike simulated datasets, DAIR-V2X introduces the first real-world dataset for cooperative detection. V2X-Seq [49] extends sequences from DAIR-V2X [48] with track IDs, establishing a sequential perception and trajectory forecasting dataset. V2V4Real [39] presents the initial real-world V2V dataset gathered from two connected vehicles. V2X-Real [32] is a large-scale multi-modal multi-view dataset intended for V2X research, compiled using two connected automated vehicles and two intelligent roadside units. TUMTraf-V2X [55] is a multimodal, multi-view V2X cooperative perception dataset designed for 3D object detection and tracking in traffic scenarios. A common limitation among the cooperative perception datasets mentioned above is their narrow focus on camera and LiDAR sensors, overlooking the benefits of 4D Radar. 4D Radar is recognized for its superior adaptability to adverse weather conditions, as demonstrated by single-vehicle autonomous driving datasets such as K-Radar [19] and Dual-Radar [52]. Consequently, It’s essential to create a comprehensive multimodal dataset that integrates 4D Radar to enhance cooperative perception research.
@@ 2.2 Cooperative Perception
Cooperative perception aims to extend the range of perception and overcome occlusions in single-vehicle perception by leveraging shared information among connected agents. It can be classified into three main types based on fusion strategies: (1) Early Fusion, which involves transmitting raw sensor data for the ego vehicle to aggregate and predict objects; (2) Late Fusion, which integrates detection results for a consistent prediction; and (3) Intermediate Fusion, which shares and fuses intermediate high-level features. Recent state-of-the-art methods [30, 2, 35, 36, 8, 18, 30, 4, 50, 51, 5] typically follow intermediate fusion, achieving the best trade-off between accuracy and bandwidth requirements. For instance, V2VNet [30] uses a graph neural network to refine features and perform joint perception and prediction iteratively. F-Cooper [2] introduces a pooling mechanism for identifying salient features. CoBEVT [35] introduced local-global sparse attention for improved cooperative BEV map segmentation. V2X-ViT [36] presented a unified vision transformer for multi-agent and multi-scale perception. HM-ViT [8] introduced a 3D heterogeneous graph transformer for camera and LiDAR fusion. FFNet [50] is a novel flow-based feature fusion framework that addresses temporal asynchrony in cooperative 3D object detection through feature flow prediction. HEAL [18] proposed an extensible collaborative perception framework to integrate new heterogeneous agents with minimal integration cost and performance decline.
@@ 3.1 Data Acquisition
Infra. LiDAR RoboSense RS-Ruby-80 ($\times$1) 80 beams, $360^{\circ}$ horizontal FOV, $-25^{\circ}\text{to}+25^{\circ}$vertical FOV
Camera Basler acA1920-40gc ($\times$3) RGB, 1536×864 resolution
4D Radar OCULI EAGLE ($\times$1) 79.0GHz, $-56^{\circ}\text{to}+56^{\circ}$ horizontal FOV, $-22^{\circ}\text{to}+22^{\circ}$ vertical FOV
@@
Vehicle LiDAR RoboSense RS-Ruby-80 ($\times$1) 80 beams, $360^{\circ}$ horizontal FOV, $-25^{\circ}\text{to}+25^{\circ}$ vertical FOV
Camera Basler acA1920-40gc ($\times$1) RGB, 1920×1080 resolution
4D Radar Arbe Phoenix ($\times$1) 77GHz, $-50^{\circ}\text{to}+50^{\circ}$ horizontal FOV, $-15^{\circ}\text{to}+15^{\circ}$ vertical FOV
@@
Figure 3: Visualization of calibration results. a) The calibration results between the camera and LiDAR. b) The calibration results between the 4D Radar and LiDAR / Camera. The LiDAR points are projected onto the camera plane using the camera’s intrinsic parameters and the camera-LiDAR extrinsics. Similarly, the 4D Radar points are transferred to the LiDAR coordinate system using the 4D Radar-LiDAR extrinsics. Additionally, these 4D Radar points are also mapped onto the camera plane by employing the camera’s intrinsic parameters, along with the extrinsic parameters of the camera-LiDAR and 4D Radar-LiDAR.
@@
Data Collection. We collected 15 hours of driving data, comprising 540K frames and capturing various weather conditions, such as sunny, rainy, foggy, and snowy, across daytime, dusk, and nighttime, as well as typical challenging intersection scenarios. Further details can be found in the supplementary materials. From this, we manually selected the most representative 40 sequences to form V2X-Radar-C; each sequence lasts between 10 to 25 seconds, with a capturing frequency of 10 Hz. Based on this, we also sampled 10 vehicle-only sequences to create V2X-Radar-V and sampled an additional 10 infrastructure-only sequences to form V2X-Radar-I. In comparison to the single-view data in V2X-Radar-C, both V2X-Radar-V and V2X-Radar-I encompass a wider variety of scenes. Consequently, these three sub-datasets collectively included 20K LiDAR frames, 40K camera images, and 20K 4D Radar data.
@@
3D Bounding Boxes Annotation. Data annotation encompasses annotations for vehicles, roadsides, and collaborative efforts. Vehicle and roadside annotations are manually created within their respective LiDAR coordinate systems. For collaborative annotation, we first align the vehicle-side and roadside annotations within a unified roadside LiDAR coordinate system, then match them using the Intersection over Union (IoU) metric to remove duplicates. The annotation process involves five object categories: pedestrian, cyclist, car, bus, and truck. Each object annotation includes attributes such as occlusion and truncation, alongside the geometric properties of 3D bounding boxes, denoted as $(x,y,z,w,h,l,\theta)$, where $(l.w.h)$ represent the object’s dimensions, $(x,y,z)$ indicate its location, and $heta$ denotes its orientation.
@@ 3.3 Data Analysis
Fig. 5(a) illustrates the distribution of objects across five categories under both day and night conditions, with cars being the most prevalent in V2X-Radar, followed by cyclists and pedestrians, while trucks and buses are the least common. Fig. 5(b) displays the maximum and average number of LiDAR points within 3D bounding boxes for each category, suggesting that larger vehicles possess more 3D points than smaller pedestrians or cyclists. Fig. 5(c) reveals the density distribution of 4D Radar points within various objects’ bounding boxes, reflecting a trend similar to that shown in Fig. 5(b). Lastly, Fig. 5(d) indicates that annotations per collaborative sample can reach up to 90 - a significant increase compared to single-vehicle datasets such as KITTI [6] or nuScenes [1], emphasising how the integration of vehicle-side and roadside data enhances the comprehensiveness of environmental perception.

### v3

@@ Abstract
Modern autonomous vehicle perception systems often struggle with occlusions and limited perception range. Previous studies have demonstrated the effectiveness of cooperative perception in extending the perception range and overcoming occlusions, thereby enhancing the safety of autonomous driving. In recent years, a series of cooperative perception datasets have emerged; however, these datasets primarily focus on cameras and LiDAR, neglecting 4D sensor used in single-vehicle autonomous driving to provide robust perception in adverse weather conditions. In this paper, to bridge the gap created by the absence of 4D Radar datasets in cooperative perception, we present V2X-Radar, the first large-scale, real-world multi-modal dataset featuring 4D Radar. V2X-Radar dataset is collected using a connected vehicle platform and an intelligent roadside unit equipped with 4D Radar, LiDAR, and multi-view cameras. The collected data encompasses sunny and rainy weather conditions, spanning daytime, dusk, and nighttime, as well as various typical challenging scenarios. The dataset consists of 20K LiDAR frames, 40K camera images, and 20K 4D Radar data, including 350K annotated boxes across five categories. To support various research domains, we have established V2X-Radar-C for cooperative perception, V2X-Radar-I for roadside perception, and V2X-Radar-V for single-vehicle perception. Furthermore, we provide comprehensive benchmarks across these three sub-datasets. We will release all datasets and benchmark codebase [^1] https://github.com/yanglei18/V2X-Radar.
footnotetext: † Corresponding author.
@@ 1 Introduction
Perception is critical in autonomous driving. While many single-vehicle perception methods [32; 31; 15; 27; 61; 47; 25; 24; 28; 14; 48; 30] have emerged, substantial safety challenges persist due to occlusions and limited perception ranges. These issues occur because vehicles can only view their surroundings from a single perspective, resulting in an incomplete understanding of the scenario. This limitation hinders autonomous vehicles from achieving safe navigation and making optimal decisions. To address this challenge, recent studies have explored cooperative perception, where ego vehicles extend their perception range and overcome occlusions with assistance from other vehicles or roadside perception [45; 44; 46; 49; 34].
@@
Figure 1: A data frame sampled from the V2X-Radar dataset. Each sample includes data from three sensors: (1) dense point clouds (gray points) from roadside and vehicle-side LiDAR; (2) sparse point clouds (green and blue points) with Doppler information from the roadside and vehicle-mounted 4D Radar; (3) RGB images (top row) from the vehicle-side camera and multi-view roadside cameras. All sensors are temporally and spatially synchronized. Each data frame is manually annotated with 3D boxes across five categories.
Table 1: Comparisons of our proposed V2X-Radar with existing V2X datasets. "C": Camera, "L": Lidar, "R": Radar, "V2V": vehicle-to-vehicle, "V2I": Vehicle-to-infrastructure.
Dataset Location Sensor V2X Real 4D Radar Adverse Weather Day&Night #LiDAR #Images #4D Radar #3D Box
OpenV2V [42] Carla L&C V2V 11K 44K 0 233K
V2X-Set [41] Carla L&C V2V&I 11K 44K 0 233K
DAIR-V2X [54] China L&C V2I 39K 39K 0 464K
V2X-Seq [56] China L&C V2I 39K 39K 0 464K
V2V4Real [40] USA L&C V2V 20K 40K 0 240K
TUMTraf-V2X[62] DE L&C V2I 2K 5K 0 29.38K
RCooper [7] China L&C I2I 30K 50K 0 310K
0.3451 0.03137 0.45882V0.39608 0.05882 0.4078420.44706 0.0902 0.35686X0.49804 0.11765 0.30588-0.55294 0.14902 0.25882R0.60392 0.17647 0.20784a0.6549 0.20784 0.15686d0.70588 0.23529 0.10588a0.75686 0.26667 0.0549r\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: China L&C&R V2I 20K 40K 20K 350K
To address the gap in the 4D Radar dataset for cooperative perception and to facilitate related studies, we present V2X-Radar, the first large-scale real-world multi-modal dataset featuring 4D Radar. This dataset includes a diverse range of scenarios, covering various weather conditions such as sunny, rainy, and snowy environments, and spans different times of the day, including daytime, dusk, and nighttime. The data was collected using a connected vehicle platform alongside an intelligent roadside unit, both equipped with 4D Radar, LiDAR, and multi-view cameras (Fig. 1). From over 15 hours of driving logs, we meticulously selected 50 representative scenarios for the final dataset. It comprises 20K LiDAR frames, 40K camera images, and 20K 4D Radar data, featuring 350K annotated bounding boxes across five object categories. To support a variety of research areas, V2X-Radar is further divided into three specialised sub-datasets: V2X-Radar-C for cooperative perception, V2X-Radar-I for roadside perception, and V2X-Radar-V for single-vehicle perception. Additionally, we provide comprehensive benchmarks of recent perception algorithms across these three sub-datasets. In comparison to existing real-world cooperative datasets, our V2X-Radar demonstrates two key strengths: (1) More modalities: The proposed V2X-Radar dataset includes three types of sensors: LiDAR, Camera, and 4D Radar, enabling further exploration into 4D Radar-related cooperative perception research. (2) Diverse scenarios: Our data collection covers diverse traffic densities, weather conditions (rain, fog, and snow), and times of day, focusing on complex intersections that pose challenges for single-vehicle autonomous driving. These scenarios include obstructed blind spots that affect vehicle safety, providing rich and varied corner cases for advancing cooperative perception research.
@@
  • Comprehensive benchmarks of recent perception algorithms are conducted across cooperative perception on V2X-Radar-C, roadside perception on V2X-Radar-I, and vehicle-side perception on V2X-Radar-V. We will release the entire V2X-Radar dataset and benchmark codebase.
@@ 2.1 Autonomous Driving Datasets
Public datasets [1; 29; 42; 41; 13; 56; 54; 40; 35; 19; 59; 60] have significantly accelerated the progress of autonomous driving in recent years. Single-vehicle datasets, such as KITTI [6], NuScenes [1], and Waymo [29], have notably furthered the development of single-vehicle perception. In terms of cooperative perception datasets, OPV2V [42] is the first dataset in this area, gathering data through co-simulation using CARLA [3] and OpenCDA [37; 38]. V2XSet [41] and V2X-Sim [13] explore V2X perception by employing synthesized data from the CARLA simulator [3]. Unlike simulated datasets, DAIR-V2X introduces the first real-world dataset for cooperative detection. V2X-Seq [56] extends sequences from DAIR-V2X [54] with track IDs, establishing a sequential perception and trajectory forecasting dataset. V2V4Real [40] presents the initial real-world V2V dataset gathered from two connected vehicles. V2X-Real [35] is a large-scale multi-modal multi-view dataset intended for V2X research, compiled using two connected automated vehicles and two intelligent roadside units. TUMTraf-V2X [62] is a multimodal, multi-view V2X cooperative perception dataset designed for 3D object detection and tracking in traffic scenarios. A common limitation among the cooperative perception datasets mentioned above is their narrow focus on camera and LiDAR sensors, overlooking the benefits of 4D Radar. 4D Radar is recognized for its superior adaptability to adverse weather conditions, as demonstrated by single-vehicle autonomous driving datasets such as K-Radar [19] and Dual-Radar [59]. Consequently, It’s essential to create a multimodal dataset that integrates 4D Radar to enhance cooperative perception research.
@@ 2.2 Cooperative Perception
Cooperative perception aims to extend the range of perception and overcome occlusions in single-vehicle perception by leveraging shared information among connected agents. It can be classified into three main types based on fusion strategies: (1) Early Fusion, which involves transmitting raw sensor data for the ego vehicle to aggregate and predict objects; (2) Late Fusion, which integrates detection results for a consistent prediction; and (3) Intermediate Fusion, which shares and fuses intermediate high-level features. Recent state-of-the-art methods [33; 2; 39; 41; 8; 17; 33; 4; 55; 57; 5] typically follow intermediate fusion, achieving the best trade-off between accuracy and bandwidth requirements. For instance, V2VNet [33] uses a graph neural network to refine features and perform joint perception and prediction iteratively. F-Cooper [2] introduces a pooling mechanism for identifying salient features. CoBEVT [39] introduced local-global sparse attention for improved cooperative BEV map segmentation. V2X-ViT [41] presented a unified vision transformer for multi-agent and multi-scale perception. HM-ViT [8] introduced a 3D heterogeneous graph transformer for camera and LiDAR fusion. FFNet [55] is a novel flow-based feature fusion framework that addresses temporal asynchrony in cooperative 3D object detection through feature flow prediction. HEAL [17] proposed an extensible collaborative perception framework to integrate new heterogeneous agents with minimal integration cost and performance decline.
@@ 3.1 Data Acquisition
Infra. LiDAR RoboSense RS-Ruby-80 ($\times$1) 80 beams, $360^{\circ}$ horizontal FOV, $25^{\circ}\text{to}+25^{\circ}$vertical FOV
[…]
4D Radar OCULI EAGLE ($\times$1) 79.0GHz, $56^{\circ}\text{to}+56^{\circ}$ horizontal FOV, $22^{\circ}\text{to}+22^{\circ}$ vertical FOV
@@
Vehicle LiDAR RoboSense RS-Ruby-80 ($\times$1) 80 beams, $360^{\circ}$ horizontal FOV, $25^{\circ}\text{to}+25^{\circ}$ vertical FOV
[…]
4D Radar Arbe Phoenix ($\times$1) 77GHz, $50^{\circ}\text{to}+50^{\circ}$ horizontal FOV, $15^{\circ}\text{to}+15^{\circ}$ vertical FOV
@@
Figure 3: Visualization of calibration results. a) The calibration results between the camera and LiDAR. b) The calibration results between the 4D Radar and LiDAR / Camera. The LiDAR points are projected onto the camera plane using the camera’s intrinsic parameters and the camera-LiDAR extrinsics. Similarly, the 4D Radar points are transferred to the LiDAR coordinate system using the 4D Radar-LiDAR extrinsics. Additionally, these 4D Radar points are also mapped onto the camera plane by employing the camera’s intrinsic parameters, along with the extrinsic parameters.
@@
Data Collection. Our data acquisition spanned over nine months, covering diverse environments such as university campuses, public roads, and closed testing parks, and ensuring comprehensive coverage across various weather conditions (sunny, rainy, foggy, and snowy) and times of day (daytime, dusk, and nighttime). In total, we collected 15 hours of driving data, comprising approximately 540K frames and encompassing numerous challenging intersection scenarios. From this dataset, we manually selected 40 representative sequences to construct V2X-Radar-C, each lasting between 10-25 seconds at a frequency of 10 Hz. Based on this, we further sampled 10 vehicle-only sequences to form V2X-Radar-V, and 10 infrastructure-only sequences to form V2X-Radar-I. Compared to the single-view configuration in V2X-Radar-C, both V2X-Radar-V and V2X-Radar-I feature a broader range of scenes. Altogether, the three subsets contain 20K LiDAR frames, 40K camera images, and 20K 4D radar samples. Further details are provided in the supplementary materials.
Privacy Protection. Before public release, the dataset undergoes a rigorous desensitization process. All privacy-sensitive elements, including road names, positioning data, road signs, license plates, and faces, are meticulously anonymized through a “model-based detection + manual verification” pipeline, in which a deep-learning model first blurs or masks sensitive regions, followed by frame-by-frame human inspection to ensure complete, compliant, and ethically responsible anonymization.
@@
3D Bounding Boxes Annotation. All data were annotated through an "auto-labeling + manual refinement" workflow: automated tools first generated initial labels, which were then carefully reviewed and corrected by human annotators, followed by multiple rounds of quality control to ensure accuracy and reliability. The annotation process encompasses vehicle-side, roadside, and collaborative annotations. Vehicle and roadside annotations were manually created within their respective LiDAR coordinate systems, while for collaborative annotation, the vehicle-side and roadside annotations were first aligned into a unified roadside LiDAR coordinate system and then matched using an IoU-based strategy to remove duplicates. Five object categories were annotated: pedestrian, cyclist, car, bus, and truck. Each object is represented as $(x,y,z,w,h,l,\theta)$, where $(l,w,h)$ represent the object’s dimensions, $(x,y,z)$ indicate its location, and $\theta$ denotes its orientation.
@@ 3.3 Data Analysis
Fig. 5(a) illustrates the distribution of objects across five categories under both day and night conditions, with cars being the most prevalent in V2X-Radar, followed by cyclists and pedestrians, while trucks and buses are the least common. Fig. 5(b) displays the maximum and average number of LiDAR points within 3D bounding boxes for each category, suggesting that larger vehicles possess more 3D points than smaller pedestrians or cyclists. Fig. 5(c) reveals the density distribution of 4D Radar points within various objects’ bounding boxes, reflecting a trend similar to that shown in Fig. 5(b). Lastly, Fig. 5(d) indicates that annotations per collaborative sample can reach up to 90, a significant increase compared to single-vehicle datasets such as KITTI [6] or nuScenes [1], emphasising how the integration of vehicle-side and roadside data enhances the comprehensiveness of environmental perception.
To better characterize the data composition, we performed a comprehensive statistical analysis on the V2X-Radar-V subset. The distribution results, shown in Fig. 6, demonstrate that the dataset covers a wide range of time-of-day conditions, including morning, afternoon, dusk, and night, as well as diverse weather scenarios such as sunny, fog, rain, and snow. Such balanced coverage captures both normal and adverse environmental conditions, providing a realistic and challenging benchmark for cooperative perception under varying illumination and visibility levels.
Figure 6: Dataset distribution by time-of-day and weather conditions. a) Distribution of data across different times of day (AM, PM, dusk, night). b) Distribution of data under various weather conditions (sunny, fog, rain, snow). The dataset spans multiple temporal and environmental scenarios, reflecting a balanced and realistic coverage of real-world driving conditions.